# Interplay of superconductivity and charge-density-wave order in kagome materials

### v1

@@ I Introduction
The discovery of CDW order and superconductivity in the AVS kagome metals stimulated a significant effort to determine the microscopic structure and origin of the ordered phases. For this purpose, one often studies charge density waves or superconductivity on the kagome lattice independently. Focusing on unconventional pairing interactions, several theoretical studies have suggested a close competition of various channels with an extended $s$-wave or a chiral $d$-wave state often the leading instability [49, 50, 51, 52]. In contrast, an (orbital-selective) phonon-mediated pairing attraction was suggested in Ref. 53. However, the interplay of orders can lead to new phenomena that cannot be understood from studying each order by itself [5]. Therefore, examining the mutual influence of individual orders can be crucial for understanding the phase diagram, and additionally provide insights into the mechanism driving each phenomenon.
@@
Previous studies including both CDW order and superconductivity have mainly focused on simplified microscopic models [54, 55, 56] or within Ginzburg-Landau theory restricted to specific charge-order and pairing symmetries [57]. Stipulating on-site attraction and thus, conventional $s$-wave pairing, a time-reversal-symmetry-breaking flux phase can induce anisotropy in the excitation gap [54, 55] and even lead to a topologically non-trivial state [56], while the system with a bond-ordered CDW retains an isotropic gap. Considering only a flux phase and superconductivity within a patch model valid close to the van Hove singularity, the interaction of $s$- or $d$-wave pairing was, further, investigated within Ginzburg-Landau theory [57]. Finally, a recent study explored the possibility of a pair-density-wave instability and the resulting orders interacting with a charge density wave [58] within a self-consistent mean-field theory. Yet, no full classification of superconducting orders and their interplay with various forms of charge order has been presented up to date.
@@
Here, we wish to comprehensively study the interplay of charge-density-wave and superconducting orders on the kagome lattice, including their potential for new states. While indications of additional order with different periodicity were reported in surface-sensitive probes [59, 60, 61], we focus on a commensurate (in-plane) CDW leading to a $2\times 2$ enlarged unit cell consistent with X-ray studies in all three compounds [62, 63, 64] and study its interplay with homogeneous, in other words total momentum ${\bm{q}}={\bm{0}}$, superconductivity and pair-density-wave order of the same wave vectors as the CDW, ${\bm{q}}={\bm{M}}_{i}$. In the spirit of the framework recently introduced in Ref. 65, we derive the Ginzburg-Landau (GL) free energy of possible CDWs and superconducting orders and explore the consequences of the dominant coupling terms close to the superconducting transition temperature. Guided by the largely different temperature scales of CDW and superconducting order, we treat the CDW as an effective field: not as a free parameter of the GL free energy to be minimized for, but rather as an input informed by experiments. Importantly, our paper clarifies the role of additional point-group or time-reversal symmetry breaking of the charge density wave for different superconducting scenarios.
@@ II.1 CDW-enlarged unit cell
To derive a GL free energy, we start by discussing the relevant symmetry of the system. Considering, for simplicity, only a two-dimensional kagome layer and further neglecting spin-orbit coupling, we use $C_{6v}$ as the point group of the kagome plane. Since we are interested in commensurate order with a $2\times 2$ enlarged unit cell, we follow the scheme of Refs. 65 and 67 and classify order parameters according to the extended point group $C_{6v}^{\prime\prime\prime}$. The extended point group contains the point group operations that leave the kagome lattice invariant and additional translations $t_{i}$, see Fig. 1, which do not map the enlarged unit cell onto itself [65], thus allowing for orders that transform non-trivially under these translations. In addition to the irreducible representations (irreps) $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ (all one-dimensional) and $E_{1}$, $E_{2}$ (two-dimensional) inherited from $C_{6v}$, the $C_{6v}^{\prime\prime\prime}$ point group contains four three-dimensional irreducible representations $F_{n}$, $n=1,\ldots,4$, which explicitly break the translational symmetry of the pristine kagome lattice  [^1] Note that the assignment of mirrors $\sigma_{v}$ and $\sigma_{d}$ relative to the lattice is not unique. The convention used for these mirrors then fixes the irreps $B_{1}$, $B_{2}$, $F_{3}$ and $F_{4}$. Furthermore, in the notation of Ref. 68, the $F_{n}$ irreps correspond to, respectively: $F_{1}\sim M_{1}^{+}$, $F_{2}\sim M_{2}^{+}$, $F_{3}\sim M_{4}^{-}$, and $F_{4}\sim M_{3}^{-}$.. The character table of the extended point group $C_{6v}^{\prime\prime\prime}$ is given in App. A.
We classify electronic orders corresponding to finite mean-field expectation values distinguishing particle-hole ($\braket{c^{\dagger}_{i\sigma}c^{\phantom{{\dagger}}}_{j\sigma^{\prime}}}$) and particle-particle ($\braket{c_{i\sigma}c_{j\sigma^{\prime}}}$) channels, where $c^{{\dagger}}_{i\sigma}$ creates an electron on site $i$ with spin $\sigma$. In particular, we ask what order-parameter symmetries can be realized given a range $r=|i-j|$. For the convenience of the later discussion, we first summarize the results of Ref. 67 on charge ordering. For on-site (OS) charge order (CO), $i=j$, the resulting mean-field is purely real and can be interpreted as a local potential. The twelve sites within the enlarged unit cell can host twelve orders, with the permutation representation $\mathcal{P}^{\rm CO}_{\rm OS}$ of these orders decomposing into $\mathcal{P}^{\rm CO}_{\rm OS}=A_{1}\oplus E_{2}\oplus F_{1}\oplus F_{3}\oplus F_{4}$. Concretely, we find three translationally invariant intra-unit-cell orders and nine CDW-type orders [65, 67].
@@
The pairing potentials corresponding to the trivial irrep $A_{1}$ are given by $\Delta_{i}=\Delta$ for all sites $i$. The $E_{2}$ irrep in the above decomposition does not break translational symmetry, but is rather modulated within the pristine kagome lattice unit cell, and has been found to appear within a recent fRG analysis [73]. The irreps denoted by $F_{n}$, in contrast, correspond to translational-symmetry-breaking superconducting orders, in other words, they describe ${\bm{q}}={\bm{M}}$ pair-density waves (PDWs). The real-space representations of the $F_{n}$ pairing potentials are shown in App. B.
@@ II.2.1 GL free energy for charge orders
Denoting the three components of a real CDW order parameter as ${\bm{\rho}}=(\rho_{1},\rho_{2},\rho_{3})^{T}$, where the subscript $i=1,2,3$ refers to momentum ${\bm{M}}_{i}$, the quadratic term has the form
@@
For concreteness, we discuss in the following four types of $F_{1}$ rCDW order parameters, each with different symmetry properties. There are two isotropic phases with $|\rho_{1}|=|\rho_{2}|=|\rho_{3}|$. These two orders, the Star-of-David ($\text{sign}(\rho_{1}\rho_{2}\rho_{3})<0$) and tri-hexagonal ($\text{sign}(\rho_{1}\rho_{2}\rho_{3})>0$) orders, are as mentioned distinguished by the third-order term in the GL free energy and preserve all $C_{6v}$ point group symmetries. Furthermore, we consider the phase represented by ${\bm{\rho}}=\pm\rho(1+\delta,1,1)^{T}$ ($\delta\neq 0$) with point group $C_{2v}$, which spontaneously breaks $C_{3}$ rotational symmetry and we refer to as nematic. Finally, the phase sometimes referred to as (structurally) chiral has $|\rho_{1}|\neq|\rho_{2}|\neq|\rho_{3}|$ and breaks all mirror symmetries of $C_{6v}$ in addition to $C_{3}$ rotational symmetry  [^3] Note that none of the orders considered break inversion, as they conserve $C_{2}$ and, in three dimensions, have an additional $z\mapsto-z$ mirror..
@@
Note that the phase discussed in Ref. 26 corresponds here to the nematic rCDW solution with a flux phase of the form ${\bm{\rho}}^{\prime}=i(0,\rho^{\prime},-\rho^{\prime})$ and has no magnetic moment.
@@ II.2.2 GL free energy for superconducting orders
When considering superconducting order parameters, the main difference to either the purely real or imaginary order parameters considered so far is that the superconducting order parameter is a complex quantity. Since the free energy does not depend on the order parameter’s phase, it has to be invariant under the associated $U(1)$ (gauge) symmetry. This puts the constraint that only an even number ($2n$) of superconducting order parameters can appear in the free energy. Thus, an order parameter and its complex conjugate can be treated as separate order parameters and the decomposition of their product (as opposed to their symmetrized product) has to contain $A_{1}$. Specifically, for an order parameter transforming according to the irrep $\Gamma$, a term in the free energy can be constructed if $A_{1}\subseteq(\otimes_{n}\Gamma)^{\rm S}\otimes(\otimes_{n}\Gamma^{*})^{\rm S}$. Based on these considerations, we construct in the next section the GL free energy for different homogeneous (${\bm{q}}={\bm{0}}$) superconducting order parameters, as well as their direct coupling to CDW orders and mutual coupling to PDW-type orders. For this purpose, we start with superconducting order parameters belonging to a one-dimensional irrep, such as (extended) $s$-wave order belonging to the $A_{1}$ irrep or spin-triplet $f$-wave order belonging to the $B_{2}$ irrep before moving to the two-dimensional irrep $E_{2}$, which can realize a (chiral) $d$-wave order parameter. In both cases, we can ask (1) what the influence of the CDW on the (homogeneous) superconductivity is and (2) what kind of modulation in the form of a PDW order parameter can be induced. In the subsequent section, we consider a PDW as the leading instability and consider its coupling to different ${\bm{q}}={\bm{0}}$ order parameters through either the CDW or directly. As previously mentioned, we treat in both parts the CDW as an effective field interacting with the superconducting order, rather than a dynamical variable to be determined through minimization of the free energy.
@@ III.1 One-dimensional irrep
where $a(T)=a_{0}(T-T_{\rm c,0})$ changes sign at the superconducting transition temperature $T_{\rm c,0}$, and $a_{0}$ and $b$ are phenomenological parameters. For concreteness, we will here focus on order parameters belonging to $A_{1}$, but comment on differences to other irreps whenever applicable.
@@
$\mathcal{F}^{(1,1,2)}[\eta,{\bm{\eta}}_{E_{2}};{\bm{\rho}}]=\frac{\mu}{2}\eta{\bm{\eta}}_{E_{2}}^{*}\cdot{\bm{\rho}}^{\rm ani.}+{\it c.c.},$ (11)
@@
$\mathcal{F}^{(2)}[{\bm{\eta}}]=a_{\rm PDW}|{\bm{\eta}}|^{2},$ (12)
$\mathcal{F}^{(1,1,1)}[\eta,{\bm{\eta}};{\bm{\rho}}]=\frac{\gamma}{2}\sum_{i}\rho_{i}(\eta\eta_{i}^{*}+{\it c.c.}).$ (13)
For the case of an $A_{1}$ order parameter, the coupling induces a PDW that transforms according to the same irrep as the CDW, in other words, an $F_{n}$ CDW induces an $F_{n}$ PDW. The induced PDWs for the other ${\bm{q}}={\bm{0}}$ irreps are summarized in Tab. 1 for an $F_{1}$ CDW. We emphasize that, although the order parameter for the modulated component resembles that of a pair density wave, the underlying mechanism differs. Here, the pair-density-wave component is induced by the coexistence of a charge density wave and homogeneous superconducting order—it represents a secondary order rather than the primary instability. Minimizing the free energy in Eqs. (12) and (13) with respect to $\eta_{i}^{*}$, we find the induced PDW
@@
$\mathcal{F}^{(1,1,2)}[\eta,{\bm{\eta}};{\bm{\rho}}]=\frac{\nu}{2}[\rho_{2}\rho_{3}(\eta\eta_{1}^{*}+{\it c.c.})+{\rm cyclic}].$ (15)
@@
and similarly for the other components. Consequently, the chirality can be reversed depending on the parameter values.
@@
$\mathcal{F}^{(1,1,1)}[\eta,{\bm{\eta}};{\bm{\rho}}^{\prime}]=i\frac{\gamma^{\prime}}{2}\sum_{i}\rho_{i}^{\prime}(\eta\eta_{i}^{*}-{\it c.c.}).$ (17)
@@
$\mathcal{F}^{(1,1,1,1)}[\eta,{\bm{\eta}};{\bm{\rho}},{\bm{\rho}}^{\prime}]$ $\propto i[(\rho_{2}\rho_{3}^{\prime}-\rho_{2}^{\prime}\rho_{3})(\eta\eta_{1}^{*}-c.c.)$ (19)

### v2

@@ I Introduction
The discovery of CDW order and superconductivity in the AVS kagome metals has stimulated a significant effort to determine the microscopic structure and origin of the ordered phases. For this purpose, one often studies charge density waves or superconductivity on the kagome lattice independently. Focusing on unconventional pairing interactions, several theoretical studies have suggested a close competition of various channels with an extended $s$-wave or a chiral $d$-wave state often the leading instability [49, 50, 51, 52, 53, 54, 55]. In contrast, an (orbital-selective) phonon-mediated pairing attraction was suggested in Ref. 56. However, the interplay of orders can lead to new phenomena that cannot be understood from studying each order by itself [5]. Therefore, examining the mutual influence of individual orders can be crucial for understanding the phase diagram, and additionally provide insights into the mechanism driving each phenomenon.
@@
Previous studies including both CDW order and superconductivity have mainly focused on simplified microscopic models [57, 58, 59] or within Ginzburg-Landau theory restricted to specific charge-order and pairing symmetries [60]. Stipulating on-site attraction and thus, conventional $s$-wave pairing, a time-reversal-symmetry-breaking flux phase can induce anisotropy in the excitation gap [57, 58] and even lead to a topologically non-trivial state [59], while the system with a bond-ordered CDW retains an isotropic gap. Considering only a flux phase and superconductivity within a patch model valid close to the van Hove singularity, the interaction of $s$- or $d$-wave pairing was, further, investigated within Ginzburg-Landau theory [60]. Finally, a recent study explored the possibility of a pair-density-wave instability and the resulting orders interacting with a charge density wave within a self-consistent mean-field theory [61]. Yet, no full classification of superconducting orders and their interplay with various forms of charge order has been presented up to date.
@@
Here, we wish to comprehensively study the interplay of charge-density-wave and superconducting orders on the kagome lattice, including their potential for new states. While indications of additional order with different periodicity were reported in surface-sensitive probes [62, 63, 64], we focus on a commensurate (in-plane) CDW leading to a $2\times 2$ enlarged unit cell consistent with X-ray studies in all three compounds [65, 66, 67] and study its interplay with homogeneous, in other words total momentum ${\bm{q}}={\bm{0}}$, superconductivity and pair-density-wave order of the same wave vectors as the CDW, ${\bm{q}}={\bm{M}}_{i}$. In the spirit of the framework recently introduced in Ref. 68, we derive the Ginzburg-Landau (GL) free energy of possible CDWs and superconducting orders and explore the consequences of the dominant coupling terms close to the superconducting transition temperature. Guided by the largely different temperature scales of CDW and superconducting order, we treat the CDW as an effective field: not as a free parameter of the GL free energy to be minimized for, but rather as an input informed by experiments. Importantly, our paper clarifies the role of additional point-group or time-reversal symmetry breaking of the charge density wave for different superconducting scenarios.
@@ II.1 CDW-enlarged unit cell
To derive a GL free energy, we start by discussing the relevant symmetry of the system. Considering, for simplicity, only a two-dimensional kagome layer and further neglecting spin-orbit coupling, we use $C_{6v}$ as the point group of the kagome plane. Since we are interested in commensurate order with a $2\times 2$ enlarged unit cell, we follow the scheme of Refs. 68 and 70 and classify order parameters according to the extended point group $C_{6v}^{\prime\prime\prime}$. The extended point group contains the point group operations that leave the kagome lattice invariant and additional translations $t_{i}$, see Fig. 1, which do not map the enlarged unit cell onto itself [68], thus allowing for orders that transform non-trivially under these translations. In addition to the irreducible representations (irreps) $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ (all one-dimensional) and $E_{1}$, $E_{2}$ (two-dimensional) inherited from $C_{6v}$, the $C_{6v}^{\prime\prime\prime}$ point group contains four three-dimensional irreducible representations $F_{n}$, $n=1,\ldots,4$, which explicitly break the translational symmetry of the pristine kagome lattice  [^1] Note that the assignment of mirrors $\sigma_{v}$ and $\sigma_{d}$ relative to the lattice is not unique. The convention used for these mirrors then fixes the irreps $B_{1}$, $B_{2}$, $F_{3}$ and $F_{4}$. Furthermore, in the notation of Ref. 71, the $F_{n}$ irreps correspond to, respectively: $F_{1}\sim M_{1}^{+}$, $F_{2}\sim M_{2}^{+}$, $F_{3}\sim M_{4}^{-}$, and $F_{4}\sim M_{3}^{-}$.. The character table of the extended point group $C_{6v}^{\prime\prime\prime}$ is given in App. A.
We classify electronic orders corresponding to finite mean-field expectation values distinguishing particle-hole ($\braket{c^{\dagger}_{i\sigma}c^{\phantom{{\dagger}}}_{j\sigma^{\prime}}}$) and particle-particle ($\braket{c_{i\sigma}c_{j\sigma^{\prime}}}$) channels, where $c^{{\dagger}}_{i\sigma}$ creates an electron on site $i$ with spin $\sigma$. In particular, we ask what order-parameter symmetries can be realized given a range $r=|i-j|$. For the convenience of the later discussion, we first summarize the results of Ref. 70 on charge ordering. For on-site (OS) charge order (CO), $i=j$, the resulting mean-field is purely real and can be interpreted as a local potential. The twelve sites within the enlarged unit cell can host twelve orders, with the permutation representation $\mathcal{P}^{\rm CO}_{\rm OS}$ of these orders decomposing into $\mathcal{P}^{\rm CO}_{\rm OS}=A_{1}\oplus E_{2}\oplus F_{1}\oplus F_{3}\oplus F_{4}$. Concretely, we find three translationally invariant intra-unit-cell orders and nine CDW-type orders [68, 70].
@@
The pairing potentials corresponding to the trivial irrep $A_{1}$ are given by $\Delta_{i}=\Delta$ for all sites $i$. The $E_{2}$ irrep in the above decomposition does not break translational symmetry, but is rather modulated within the pristine kagome lattice unit cell, and has been found to appear within a recent fRG analysis [76]. The irreps denoted by $F_{n}$, in contrast, correspond to translational-symmetry-breaking superconducting orders, in other words, they describe ${\bm{q}}={\bm{M}}$ pair-density waves (PDWs). The real-space representations of the $A_{1}$, $E_{2}$ and $F_{1}$ pairing potentials are shown in Fig. 2. A complete list of real-space representations is presented in App. B.
Figure 2: Real-space representations of the on-site $A_{1}$, $E_{2}$ and $F_{1}$ spin-singlet pairing potentials. The color and its intensity at each site represent relative signs and strengths of the pairing potentials.
@@ II.2.1 GL free energy for charge orders
We denote the three components of a real CDW order parameter as ${\bm{\rho}}=(\rho_{1},\rho_{2},\rho_{3})^{T}$, where the subscript $i=1,2,3$ refers to momentum ${\bm{M}}_{i}$. The quadratic term, which, as is common, is assumed to be the only term with a temperature dependence [78], has the form
@@
For concreteness, we discuss in the following four types of $F_{1}$ rCDW order parameters, each with different symmetry properties. There are two isotropic phases with $|\rho_{1}|=|\rho_{2}|=|\rho_{3}|$. These two orders, the Star-of-David [$\text{sign}(\rho_{1}\rho_{2}\rho_{3})<0$] and tri-hexagonal [$\text{sign}(\rho_{1}\rho_{2}\rho_{3})>0$] orders, are as mentioned distinguished by the third-order term in the GL free energy and preserve all $C_{6v}$ point group symmetries. Furthermore, we consider the phase represented by ${\bm{\rho}}=\pm(\rho,\rho+\delta,\rho)^{T}$ ($\delta\neq 0$) with point group $C_{2v}$, which spontaneously breaks $C_{3}$ rotational symmetry and we refer to as nematic. Finally, the phase sometimes referred to as (structurally) chiral has $|\rho_{1}|\neq|\rho_{2}|\neq|\rho_{3}|$ and breaks all mirror symmetries of $C_{6v}$ in addition to $C_{3}$ rotational symmetry  [^3] Note that none of the orders considered break inversion, as they conserve $C_{2}$ and, in three dimensions, have an additional $z\mapsto-z$ mirror..
@@
Note that the phase discussed in Ref. 26 corresponds here to the nematic rCDW solution with a flux phase of the form ${\bm{\rho}}^{\prime}=i(\rho^{\prime},0,-\rho^{\prime})$ and has no magnetic moment.
@@ II.2.2 GL free energy for superconducting orders
When considering superconducting order parameters, the main difference to either the purely real or imaginary order parameters considered so far is that the superconducting order parameter is a complex quantity. Since the free energy has to be invariant under the $U(1)$ (gauge) symmetry, it cannot depend on the order parameter’s phase. This puts the constraint that only an even number ($2n$) of superconducting order parameters can appear in the free energy. Further, an order parameter and its complex conjugate can be treated as separate order parameters and the decomposition of their product (as opposed to their symmetrized product) has to contain $A_{1}$. Specifically, for an order parameter transforming according to the irrep $\Gamma$, a term in the free energy can be constructed if $A_{1}$ is in the decomposition of $(\otimes_{n}\Gamma)^{\rm S}\otimes(\otimes_{n}\Gamma^{*})^{\rm S}$. Based on these considerations, we construct in the next section the GL free energy for different homogeneous (${\bm{q}}={\bm{0}}$) superconducting order parameters, as well as their direct coupling to CDW orders and mutual coupling to PDW-type orders. For this purpose, we start with superconducting order parameters belonging to a one-dimensional irrep, such as (extended) $s$-wave order belonging to the $A_{1}$ irrep or spin-triplet $f$-wave order belonging to the $B_{2}$ irrep, before moving to the two-dimensional irrep $E_{2}$, which can realize a (chiral) $d$-wave order parameter. In both cases, we can ask (1) what the influence of the CDW on the (homogeneous) superconductivity is and (2) what kind of modulation in the form of a PDW order parameter can be induced, see illustration in Fig. 3. In the subsequent section, we consider a PDW as the leading instability and consider its coupling to different ${\bm{q}}={\bm{0}}$ order parameters through either the CDW or directly. As previously mentioned, we treat in both parts the CDW as an effective field interacting with the superconducting order, rather than a dynamical variable to be determined through minimization of the free energy.
@@ III Dominant homogeneous superconducting order
Figure 3: Effect of a charge density wave $\delta\rho(r)$ on a homogeneous superconducting order parameter $\eta_{0}$: (1) direct coupling changes $T_{\rm c}$ and the absolute value of $\eta_{0}$ and (2) the order parameter is generically modulated with the same wave vector as the CDW.
@@ III.1 One-dimensional irrep
where $a(T)=a_{0}(T-T_{\rm c,0})$ changes sign at the superconducting transition temperature $T_{\rm c,0}$, and $a_{0}$ and $b$ are phenomenological parameters. Note that here and in the following, only the quadratic term of the dominant superconducting order parameter is assumed to have a temperature dependence, while all the other phenomenological parameters, including the quadratic term of the secondary orders, are temperature independent. For concreteness, we will here focus on order parameters belonging to $A_{1}$, but comment on differences to other irreps whenever applicable.
@@
$\mathcal{F}^{(1,1;2)}[\eta,{\bm{\eta}}_{E_{2}};{\bm{\rho}}]=\frac{\mu}{2}\eta{\bm{\eta}}_{E_{2}}^{*}\cdot{\bm{\rho}}^{\rm ani.}+{\it c.c.},$ (11)
@@
For the case of an $A_{1}$ order parameter, the coupling induces a PDW that transforms according to the same irrep as the CDW, in other words, an $F_{n}$ CDW induces an $F_{n}$ PDW. The induced PDWs for the other ${\bm{q}}={\bm{0}}$ irreps are summarized in Tab. 1 for an $F_{1}$ CDW. We emphasize that, although the order parameter for the modulated component resembles that of a pair density wave, the underlying mechanism differs. In the case of a primary homogeneous ($\bm{q}=\bm{0}$) superconducting order, the pair-density-wave component is induced by the spatial modulation due to the charge density wave—it represents a secondary order rather than the primary instability. Minimizing the free energy in Eqs. (12) and (13) with respect to $\eta_{i}^{*}$, we find the induced PDW
@@
$\mathcal{F}^{(1,1;2)}[\eta,{\bm{\eta}};{\bm{\rho}}]=\frac{\nu}{2}[\rho_{2}\rho_{3}(\eta\eta_{1}^{*}+{\it c.c.})+{\rm cyclic}].$ (15)
@@
and similarly for the other components. Consequently, the chirality can be reversed depending on the parameter values. Figure 4 summarizes the effect of a CDW on homogeneous superconductivity, showing the amplitude of induced orders as a function of CDW anisotropy $\delta$.
Figure 4: (a) Illustration of the induced orders from a homogeneous superconducting order $|\eta|$ coexisting with a charge density wave $\bm{\rho}\propto(\rho,\rho+\delta,\rho)^{T}$ with $|\bm{\rho}|=1$. The inset illustrates the tri-hexagonal CDW order that becomes anisotropic as $\delta$ increases. Figure obtained with $a(T)=-0.25$, $a_{E_{2},0}=a_{\rm PDW,0}=b=1$, $\gamma=-0.75$, and $\mu=-0.25$. (b) Impact of the anisotropic CDW considered in (a) on the $A_{1}$ superconducting order: (1) inducing an $E_{2}$ component and (2) splitting the strengths of the induced PDW components.
@@
$\mathcal{F}^{(1,1;1)}[\eta,{\bm{\eta}};{\bm{\rho}}^{\prime}]=i\frac{\gamma^{\prime}}{2}\sum_{i}\rho_{i}^{\prime}(\eta\eta_{i}^{*}-{\it c.c.}).$ (17)
@@
$\mathcal{F}^{(1,1;1,1)}[\eta,{\bm{\eta}};{\bm{\rho}},{\bm{\rho}}^{\prime}]$ $\propto i[(\rho_{2}\rho_{3}^{\prime}-\rho_{2}^{\prime}\rho_{3})(\eta\eta_{1}^{*}-c.c.)$ (19)